Octane (C8H18) is a component of gasoline that burns according to the following equation:
C8H18(l)+252O2(g)→8CO2(g)+9H2O(g)
ΔH∘rxn=−5074.1kJ

What mass of octane (in g ) is required to produce 1850 kJ of heat?        
Answer #1

from reaction,

1 mol of octane produces 5074.1 KJ

So,

to produce 1850 KJ, mol of octane required = 1850/5074.1

= 0.3646 mol

Molar mass of C8H18 = 8*MM(C) + 18*MM(H)

= 8*12.01 + 18*1.008

= 114.224 g/mol

we have below equation to be used:

mass of C8H18,

m = number of mol * molar mass

= 0.3646 mol * 114.224 g/mol

= 41.7 g

Answer: 41.7 g

Feel free to comment below if you have any doubts or if this answer do not work